OpenStudy (anonymous):

A football team lost 9 yards on its first play. On the next two plays they gained 3 yards and then 6 yards. write an expression to show this

OpenStudy (anonymous):

@Learner11

OpenStudy (anonymous):

@texaschic101

OpenStudy (nykoreyan):

9-9=0

OpenStudy (anonymous):

lol good job

OpenStudy (anonymous):

I was just giving a free medal to someone

OpenStudy (nykoreyan):

lmao

OpenStudy (anonymous):

-9+3+3+6 = 12-9 = 3 @Shadowknight6263 , Sorry for the late reply :(

OpenStudy (anonymous):

it's fine :)

OpenStudy (anonymous):

:)